# Channel-Imposed Fusion: A Simple yet Effective Method for Medical Time Series Classification

### v1

@@ 5.3 Ablation Study
(1) Effectiveness of CIF: Table 3 demonstrates the excellent performance of combining HM-BiTCN with CIF, confirming the compatibility of the HM-BiTCN with CIF. See Appendix F.1 for details.
@@
(2) Hyperparameter Transfer and Adaptation: We evaluate the transferability of key hyperparameters (e.g., , , ) from HM-BiTCN to other models. If transferred settings underperform, we further fine-tune them for adaptation. See Appendices F.2 and F.3. Figure 4 illustrates the outstanding performance of CIF when combined with other models.
@@ 6 Conclusion
In this paper, we propose a simple yet effective temporal classification representation method called Channel-Imposed Fusion (CIF) and combine it with a straightforward TCN structure. This approach outperforms current SOTA methods on multiple medical time-series datasets. Additionally, the CIF method demonstrates strong adaptability, enabling integration with cutting-edge methods to further enhance their performance. Through the CIF method, we aim to show researchers in the field that, rather than solely modifying models to process high-level features, exploring the underlying core features of the data itself is also a research path worth delving into.
Appendix A Proof of SNR Optimization via CIF.
Consider two observed signals:
where:
are zero-mean useful signals with variance and correlation coefficient
are zero-mean noise components with variance and correlation coefficient
Signal and noise components are mutually uncorrelated
Derivation of Output SNR
For the linear combination :
@@ 2. Noise Power Calculation
3. SNR Ratio
The output SNR becomes:
@@ 3. SNR Ratio
Optimization Conditions
SNR improvement occurs when:
Difference Mode ():
Cooperative Mode ():
Appendix B Visualization of the Channel-Imposed Fusion (CIF) Process.
In Figure 5 6, we visualize scenarios with highly correlated channel information and those with generally correlated channel information. These visualizations intuitively demonstrate the strong representation capability of our proposed method.
Appendix C Data Preprocessing
| Datasets | #-Subject | #-Sample | #-Class | #-Channel | #-Timestamps | Sampling Rate | Modality | File Size |
| APAVA | 23 | 5,967 | 2 | 16 | 256 | 256Hz | EEG | 186MB |
| ADFTD | 88 | 69,752 | 3 | 19 | 256 | 256Hz | EEG | 2.52GB |
| TDBrain | 72 | 6,240 | 2 | 33 | 256 | 256Hz | EEG | 571MB |
| PTB | 198 | 64,356 | 2 | 15 | 300 | 250Hz | ECG | 2.15GB |
| PTB-XL | 17,596 | 191,400 | 5 | 12 | 250 | 250Hz | ECG | 4.28GB |
Appendix D Implementation Details
We implement our method and all the baselines based on the Time-Series-Library project111https://github.com/thuml/Time-Series-Library from Tsinghua University Wu et al. [2023], which integrates all methods under the same framework and training techniques to ensure a relatively fair comparison. The 12 baseline time series transformer methods are Autoformer Wu et al. [2021], Crossformer Zhang and Yan [2022], FEDformer Zhou et al. [2022], Informer Zhou et al. [2021], iTransformer Liu et al. [2024b], MTST Zhang et al. [2024], Nonformer Liu et al. [2022], PatchTST Nie et al. [2023], Reformer Kitaev et al. [2019], vanilla Transformer Vaswani et al. [2017], Medformer Wang et al. [2024], and MedGNN Fan et al. [2025].
Autoformer Autoformer Wu et al. [2021] employs an auto-correlation mechanism to replace self-attention for time series forecasting. Additionally, they use a time series decomposition block to separate the time series into trend-cyclical and seasonal components for improved learning. The raw source code is available at https://github.com/thuml/Autoformer.
Crossformer Crossformer Zhang and Yan [2022] designs a single-channel patching approach for token embedding. They utilize two-stage self-attention to leverage both temporal features and channel correlations. A router mechanism is proposed to reduce time and space complexity during the cross-dimension stage. The raw code is available at https://github.com/Thinklab-SJTU/Crossformer.
FEDformer FEDformer Zhou et al. [2022] leverages frequency domain information using the Fourier transform. They introduce frequency-enhanced blocks and frequency-enhanced attention, which are computed in the frequency domain. A novel time series decomposition method replaces the layer norm module in the transformer architecture to improve learning. The raw code is available at https://github.com/MAZiqing/FEDformer.
Informer Informer Zhou et al. [2021] is the first paper to employ a one-forward procedure instead of an autoregressive method in time series forecasting tasks. They introduce ProbSparse self-attention to reduce complexity and memory usage. The raw code is available at https://github.com/zhouhaoyi/Informer2020.
iTransformer iTransformer Liu et al. [2024b] questions the conventional approach of embedding attention tokens in time series forecasting tasks and proposes an inverted approach by embedding the whole series of channels into a token. They also invert the dimension of other transformer modules, such as the layer norm and feed-forward networks. The raw code is available at https://github.com/thuml/iTransformer.
MTST MTST Zhang et al. [2024] uses the same token embedding method as Crossformer and PatchTST. It highlights the importance of different patching lengths in forecasting tasks and designs a method that can take different sizes of patch tokens as input simultaneously. The raw code is available at https://github.com/networkslab/MTST.
Nonformer Nonformer Liu et al. [2022] analyzes the impact of non-stationarity in time series forecasting tasks and its significant effect on results. They design a de-stationary attention module and incorporate normalization and denormalization steps before and after training to alleviate the over-stationarization problem. The raw code is available at https://github.com/thuml/Nonstationary_Transformers.
PatchTST PatchTST Nie et al. [2023] embeds a sequence of single-channel timestamps as a patch token to replace the attention token used in the vanilla transformer. This approach enlarges the receptive field and enhances forecasting ability. The raw code is available at https://github.com/yuqinie98/PatchTST.
Reformer Reformer Kitaev et al. [2019] replaces dot-product attention with locality-sensitive hashing. They also use a reversible residual layer instead of standard residuals. The raw code is available at https://github.com/lucidrains/reformer-pytorch.
Transformer Transformer Vaswani et al. [2017], commonly known as the vanilla transformer, is introduced in the well-known paper "Attention is All You Need." It can also be applied to time series by embedding each timestamp of all channels as an attention token. The PyTorch version of the code is available at https://github.com/jadore801120/attention-is-all-you-need-pytorch.
Medformer Medformer Wang et al. [2024] uses cross-channel patch embedding to model spatiotemporal dependencies.The raw code is available at https://github.com/DL4mHealth/Medformer
MedGNN MedGNN Fan et al. [2025] employs multi-resolution spatiotemporal graph learning to extract dynamic features across multiple time scales. The raw code is available at https://github.com/aikunyi/MedGNN.
D.1 Evaluation metrics
@@
Appendix E Reproduced Results of Existing Methods
We employ six evaluation metrics: accuracy, precision (macro-averaged), recall (macro-averaged), F1 score (macro-averaged), AUROC (macro-averaged), and AUPRC (macro-averaged). The training process is conducted with five random seeds (41-45) on fixed training, validation, and test sets to compute the mean and standard deviation of the models. All experiments are run on an NVIDIA RTX 3090 GPU.
E.1 Results of Subject-Dependent
In Table 5, we present the results of various models under the Subject-Dependent Setup on the ADFTD (3-Classes) dataset. The results indicate that our proposed CIF method significantly improves the performance of models, particularly in accuracy, F1 score, AUROC, and AUPRC metrics. Specifically, the combination of MedGNN + CIF outperforms MedGNN alone across all key performance metrics, and similarly, Medformer + CIF surpasses Medformer in every critical evaluation metric. These results convincingly demonstrate the effectiveness of the CIF method in enhancing model performance. By integrating the CIF method, not only are the models’ performances significantly improved across multiple evaluation metrics, but the broad applicability and substantial benefits of the method across different model architectures are also highlighted. Notably, MedGNN + CIF achieves a remarkable Accuracy of 99.60%, far surpassing other models, highlighting the advantage of our approach in handling imbalanced datasets. Furthermore, MedGNN + CIF shows significant improvements in F1 score and precision, indicating that the CIF method effectively enhances the model’s ability to recognize both positive and negative samples. Thus, our experimental results validate the effectiveness of the CIF method in boosting model generalization and precision.
@@ E.1 Results of Subject-Dependent
| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
| ADFTD (3-Classes) Reproduced | Autoformer Wu et al. [2021] | 87.792.87 | 87.632.95 | 86.843.10 | 87.173.05 | 96.451.15 | 93.522.05 |
| Crossformer Zhang and Yan [2022] | 89.271.19 | 88.981.27 | 88.611.35 | 88.771.29 | 97.530.48 | 95.470.87 | |
| FEDformer Zhou et al. [2022] | 77.552.42 | 76.712.14 | 76.462.52 | 76.442.52 | 91.661.36 | 84.992.15 | |
| Informer Zhou et al. [2021] | 91.720.13 | 91.500.16 | 91.360.21 | 91.420.14 | 98.400.07 | 96.920.11 | |
| iTransformer Liu et al. [2024b] | 64.600.77 | 62.220.64 | 61.950.35 | 61.870.41 | 81.460.29 | 68.810.39 | |
| MTST Zhang et al. [2024] | 65.430.79 | 64.470.65 | 63.490.55 | 63.710.55 | 81.670.51 | 69.740.69 | |
| Nonformer Liu et al. [2022] | 96.000.72 | 95.720.80 | 95.900.69 | 95.810.75 | 99.560.13 | 99.020.23 | |
| PatchTST Nie et al. [2023] | 66.550.48 | 65.610.60 | 65.260.78 | 65.260.60 | 83.270.33 | 71.960.44 | |
| Reformer Kitaev et al. [2019] | 90.742.18 | 90.872.15 | 90.612.01 | 90.442.25 | 98.690.50 | 97.570.85 | |
| Transformer Vaswani et al. [2017] | 96.620.54 | 96.460.54 | 96.450.59 | 96.450.56 | 99.690.08 | 99.310.13 | |
| Medformer Wang et al. [2024] | 95.761.03 | 95.541.10 | 95.621.05 | 95.581.08 | 99.530.18 | 99.090.36 | |
| Medformer + CIF | 98.870.26 | 98.770.27 | 98.860.27 | 98.810.27 | 99.960.01 | 99.920.03 | |
| MedGNN Fan et al. [2025] | 99.440.21 | 99.430.21 | 99.410.22 | 99.420.21 | 99.980.01 | 99.970.01 | |
E.2 Results of Subject-Independent
Table 2 presents the results of various methods reported under the subject-independent setup, while Table 6 shows the results of our reproduction of these methods. Our approach achieves the highest scores on six metrics across four out of the five datasets. For the PTB-XL dataset, compared to the results reported by other methods, our method achieves the highest AUROC, with Accuracy ranking second. In comparison with our reproduced results, our method ranks first in Accuracy, AUROC, and AUPRC, and second in Precision and Recall. Additionally, it is noteworthy that under the subject-independent setup, the F1 score for ADFTD is 55.42%, significantly lower than the 99.59% F1 score achieved under the subject-dependent setup. This comparison highlights the challenge of the subject-independent setup, which better simulates real-world scenarios.

### v2

@@ 4.3 Ablation Study
(1) Effectiveness of CIF: Table 3 demonstrates the excellent performance of combining HM-BiTCN with CIF, confirming the compatibility of the HM-BiTCN with CIF. Appendix D presents ablation studies on the HM-BiTCN architecture and the performance benefits of integrating CIF into its components.
@@
(2) Hyperparameter Transfer and Adaptation: We evaluate the transferability of key hyperparameters (e.g., , , ) from HM-BiTCN to other models. If transferred settings underperform, we further fine-tune them for adaptation. Figure 5 illustrates the outstanding performance of CIF when combined with other models.
(3) Exploring Physiologically-Informed CIF: In Appendix E, we further conducted experiments on CIF from the perspective of emphasizing more prominent biological features. The results indicate that incorporating specific biological characteristics can further enhance classification performance.
(4)Results on general time series classification tasks
Accuracy | 75.840.52 | 71.651.27 | 74.961.25 | 66.481.66 | 74.800.96 | 64.872.78 | 76.440.64 | 76.080.81 | 76.821.32 | |||||||||||||||||||
| F1 Score | 75.520.66 | 71.141.45 | 74.491.39 | 65.291.74 | 74.350.85 | 64.142.70 | 76.250.65 | 75.540.94 | 76.391.18 | ||||||||||||||||||||
Accuracy | 89.741.08 | 88.442.02 | 88.861.65 | 93.080.95 | 91.441.01 | 87.781.10 | 91.650.74 | 93.720.73 | 93.780.32 | |||||||||||||||||||
| F1 Score | 89.701.10 | 88.341.98 | 88.801.67 | 93.190.88 | 91.470.98 | 87.721.10 | 91.610.75 | 93.690.76 | 93.740.34 | ||||||||||||||||||||
To evaluate the performance of our method on general time series, we follow the design of Medformer Wang et al. (2024a) and test it on two human activity recognition (HAR) datasets: FLAAP(13,123 samples, 10 classes) Kumar & Suresh (2022) and UCI-HAR(10,299 samples, 6 classes) Anguita et al. (2013). Additionally, to conduct a more comprehensive evaluation, following TimeMixer++ Wang et al. (2025a), we used 10 multivariate datasets from the UEA Time Series Classification Archive (2018) for the assessment of classification tasks.
As shown in Table 4, and Fig. 6, the combination of HM-BiTCN and CIF consistently outperforms other architectures in general time series classification, achieving a 5.5% improvement over the original TCN and surpassing current SOTA methods. Although CIF was originally designed for MedTS, its integration with HM-BiTCN significantly outperforms Transformer-based models in both medical and general time series classification tasks, demonstrating the effectiveness of our data-centric approach.
The results in Tables 4, 7, 3 and Figure 5 show that CIF achieves significant improvements in MedTS classification, while the gains on non-medical data are relatively limited. This observation further demonstrates that a data-driven perspective is particularly effective for MedTS classification with physiological characteristics.
@@ 5 Conclusion
In this work, we propose a simple and effective method for medical time series classification, Channel-Imposed Fusion (CIF), which explicitly encodes physiological causal relationships between channels in the feature representations while enhancing the SNR of the original signals. Combined with the simple HM-BiTCN architecture, CIF surpasses existing SOTA methods on multiple medical datasets and performs strongly on general time series classification tasks, demonstrating that data-centric design enables simple models to outperform more complex architectures. More importantly, CIF exemplifies the shift from the traditional model-centric paradigm to a data-centric perspective, where structured representations grounded in physiological priors are both efficient and scalable for medical time series classification. CIF also exhibits strong transferability and can be seamlessly integrated into mainstream models such as Transformers, enhancing their applicability in medical scenarios. We hope this work encourages the community to reconsider the core of medical time series classification: should it be driven primarily by data-centric strategies or by model-centric design or both?
Appendix
Appendix A Explanation of SNR Optimization via CIF.
Consider the linear combination of the two observed signals:
where and are real coefficients. The observed signals are given by
with zero-mean signal and noise components:
and mutually uncorrelated signal and noise components: .
@@ 1. Signal Power Calculation
The power of a zero-mean random signal is given by its variance:
This is why, for zero-mean signals, the SNR can be expressed as a ratio of variances (or mean-square values) Kay (1993).
For the linear combination of signals:
where .
@@ 2. Noise Power Calculation
Similarly, the noise power of the linear combination is:
@@
3. Output SNR
Using the definition of SNR as the ratio of signal power to noise power:
where .
4. SNR Improvement Condition
For SNR improvement relative to individual channels:
Difference Mode (): — suppress correlated noise while possibly attenuating some correlated signal.
Cooperative Mode (): — amplify correlated signals relative to less-correlated noise.
A.1 Evaluation metrics
@@
The Area Under the Receiver Operating Characteristic Curve (AUROC) measures the ability of a model to distinguish between classes, defined as
where
The Area Under the Precision–Recall Curve (AUPRC) summarizes the trade-off between precision and recall across different thresholds, defined as
where
Appendix B HM-BiTCN Structure Design and Theoretical Analysis
Modeling short-term and long-term dependencies in time series data is challenging. Traditional CNNs excel at capturing local features but have limited receptive fields, hindering long-range dependency learning. Transformer-based methods effectively model long-term dependencies, but their complex design lacks interpretability, which is a key issue in medical time-series classification.To address these limitations, TCNs use causal convolutions for explicit temporal modeling and dilated convolutions to expand the receptive field, overcoming the constraints of traditional CNNs. Building on the advantages of TCN, we propose the HM-BiTCN, which combines the benefits of dilated convolutions, bidirectional causal convolution, and residual connections. This approach allows for better capture of temporal dependencies while preserving causality.
B.1 Dilated Convolution
Dilated convolution expands the receptive field without significantly increasing computational cost Yu & Koltun (2015). For a 1D input sequence , its output is defined as , where is the current time step, is the kernel size, is the dilation factor, and is the weight at the -th position in the kernel. Increasing effectively enlarges the receptive field, enabling the network to capture longer-term temporal dependencies. When stacking multiple dilated convolutional layers, the receptive field grows progressively. For the -th layer, the receptive field can be expressed as , where is the dilation factor of the -th layer. By gradually increasing , the network captures temporal dependencies across both global and local scales, offering an effective way to model long-term dependencies in time series.
B.2 Bidirectional Causal Convolution Structure
B.3 Multi-Scale Feature Learning and Residual Connections
To further improve the model’s capacity to capture dependencies at different temporal scales, HM-BiTCN incorporates multi-scale feature learning and residual connections. Multi-scale Feature Learning: In HM-BiTCN, we employ a hierarchy of dilation factors that decrease layer by layer to capture temporal dependencies at multiple scales. Lower layers use larger dilation factors to expand the receptive field, aggregating long‐range information and smoothing short‐term noise in highly redundant medical time series; higher layers use smaller dilation factors to focus on local dependencies and capture fine‐grained features. This coarse‐to‐fine, global‐to‐local design enables the network to extract broad patterns in its initial layers and refine precise details in its later layers, thereby enhancing adaptability across a wide range of time series tasks. Residual connections: Residual connections He et al. (2016) are introduced between the dilated convolutional layers to facilitate the efficient flow of information through the network. The residual connection is defined as , where is the convolutional output, and is the input. This design alleviates the vanishing gradient problem and improves the overall stability of the network during training.
Appendix C Pseudocode of CIF Method and Key Components of HM-BiTCN
Appendix D Ablation Experiments of the HM-BiTCN Structure
@@
| Datasets | Models | CIF | Forward | Backward | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC | |||
HM-BiTCN | ✓ | 82.312.34 | 83.292.50 | 80.392.65 | 81.022.63 | 91.501.80 | 91.661.82 | |||||
| HM-BiTCN | ✓ | 79.453.51 | 80.693.04 | 77.144.47 | 77.584.75 | 87.953.82 | 88.413.74 | ||||||
| HM-BiTCN | ✓ | ✓ | 82.491.40 | 82.381.79 | 81.201.32 | 81.601.39 | 91.101.63 | 91.301.71 | |||||
HM-BiTCN | ✓ | ✓ | 80.435.60 | 80.465.23 | 79.565.98 | 79.505.97 | 89.234.44 | 89.624.25 | ||||
| HM-BiTCN | ✓ | ✓ | 79.393.44 | 79.493.76 | 78.092.94 | 78.353.33 | 87.623.09 | 88.112.91 | |||||
| HM-BiTCN | ✓ | ✓ | ✓ | 85.161.55 | 84.761.62 | 85.331.27 | 84.821.49 | 94.061.07 | 94.210.99 | ||||
HM-BiTCN | ✓ | 53.321.35 | 52.011.54 | 51.462.19 | 51.211.99 | 70.781.78 | 53.162.20 | |||||
| HM-BiTCN | ✓ | 52.801.18 | 50.160.77 | 49.231.22 | 49.241.02 | 68.650.71 | 49.950.97 | ||||||
| HM-BiTCN | ✓ | ✓ | 52.052.22 | 50.453.00 | 50.402.55 | 49.482.70 | 69.432.84 | 50.993.15 | |||||
HM-BiTCN | ✓ | ✓ | 56.060.47 | 53.211.03 | 53.541.36 | 52.821.33 | 72.930.88 | 55.711.03 | ||||
| HM-BiTCN | ✓ | ✓ | 56.541.33 | 54.280.96 | 54.631.06 | 53.911.11 | 73.461.17 | 56.121.61 | |||||
| HM-BiTCN | ✓ | ✓ | ✓ | 58.560.93 | 55.650.81 | 55.860.79 | 55.420.82 | 76.070.59 | 59.750.67 | ||||
HM-BiTCN | ✓ | 87.232.87 | 87.752.48 | 87.232.87 | 87.172.93 | 95.551.69 | 95.731.60 | |||||
| HM-BiTCN | ✓ | 86.923.46 | 87.413.17 | 86.923.46 | 86.863.51 | 95.281.78 | 95.421.70 | ||||||
| HM-BiTCN | ✓ | ✓ | 84.902.60 | 86.022.00 | 84.902.60 | 84.762.74 | 93.941.92 | 94.201.85 | |||||
HM-BiTCN | ✓ | ✓ | 93.291.73 | 93.341.73 | 93.291.73 | 93.291.73 | 98.500.63 | 98.560.60 | ||||
| HM-BiTCN | ✓ | ✓ | 93.691.52 | 93.831.42 | 93.691.52 | 93.681.53 | 98.560.67 | 98.590.64 | |||||
| HM-BiTCN | ✓ | ✓ | ✓ | 93.131.41 | 93.331.37 | 93.131.41 | 93.121.42 | 98.620.66 | 98.680.63 | ||||
HM-BiTCN | ✓ | 82.561.74 | 86.161.51 | 74.912.88 | 77.242.92 | 95.690.64 | 94.560.76 | |||||
| HM-BiTCN | ✓ | 81.074.24 | 85.362.71 | 72.506.59 | 74.336.71 | 92.832.38 | 91.282.79 | ||||||
| HM-BiTCN | ✓ | ✓ | 81.871.87 | 86.501.24 | 73.492.90 | 75.843.20 | 94.200.29 | 93.040.45 | |||||
HM-BiTCN | ✓ | ✓ | 87.331.41 | 90.261.24 | 81.642.04 | 84.191.97 | 96.211.30 | 95.671.52 | ||||
| HM-BiTCN | ✓ | ✓ | 84.352.28 | 87.422.07 | 77.543.30 | 79.983.41 | 91.251.92 | 90.422.28 | |||||
| HM-BiTCN | ✓ | ✓ | ✓ | 88.291.45 | 90.661.48 | 83.212.02 | 85.591.96 | 94.280.93 | 93.781.11 | ||||
HM-BiTCN | ✓ | 70.812.31 | 72.581.33 | 69.812.79 | 70.072.24 | 95.890.28 | 76.901.21 | |||||
| HM-BiTCN | ✓ | 70.292.04 | 72.772.09 | 68.862.39 | 69.561.98 | 95.610.26 | 76.561.56 | ||||||
| HM-BiTCN | ✓ | ✓ | 76.080.81 | 76.050.83 | 75.950.84 | 75.540.94 | 96.490.10 | 81.190.65 | |||||
HM-BiTCN | ✓ | ✓ | 72.301.50 | 72.981.61 | 71.651.35 | 71.541.50 | 95.920.55 | 77.872.27 | ||||
| HM-BiTCN | ✓ | ✓ | 72.811.04 | 74.050.80 | 71.861.25 | 72.121.17 | 96.200.21 | 79.221.25 | |||||
| HM-BiTCN | ✓ | ✓ | ✓ | 76.821.32 | 77.380.85 | 76.521.24 | 76.391.18 | 96.480.06 | 81.770.81 | ||||
HM-BiTCN | ✓ | 91.940.98 | 92.360.90 | 92.020.96 | 91.980.93 | 99.300.08 | 97.310.47 | |||||
| HM-BiTCN | ✓ | 93.030.62 | 93.280.63 | 93.120.60 | 93.050.62 | 99.360.19 | 97.720.46 | ||||||
| HM-BiTCN | ✓ | ✓ | 93.720.73 | 94.020.72 | 93.750.70 | 93.690.76 | 99.600.09 | 98.310.40 | |||||
HM-BiTCN | ✓ | ✓ | 92.180.45 | 92.420.47 | 92.210.44 | 92.140.44 | 99.170.11 | 97.040.15 | ||||
| HM-BiTCN | ✓ | ✓ | 92.620.90 | 92.880.86 | 92.680.88 | 92.630.89 | 99.250.18 | 97.150.57 | |||||
| HM-BiTCN | ✓ | ✓ | ✓ | 93.780.32 | 94.080.26 | 93.790.32 | 93.740.34 | 99.340.19 | 97.600.46 |
From the table 5, it can be observed that when both the Forward and Backward parts of the HM-BiTCN structure are used simultaneously, the performance drops significantly compared to using only one of them individually. We speculate that this is mainly due to the presence of substantial noise within medical time-series data. When both parts of the structure are applied at the same time, it is akin to capturing noise from two different directions simultaneously. Instead of enhancing the representation, this leads to noise accumulation, which ultimately results in degraded performance.
However, after processing the data with CIF to improve the signal-to-noise ratio, the combination of the Forward and Backward parts of the HM-BiTCN structure eventually outperforms the use of either part alone. This result strongly demonstrates the feature-capturing capability of the HM-BiTCN when both directions are utilized together. It indicates that once noise interference is effectively reduced, the bidirectional structure of HM-BiTCN can better leverage its strengths, thereby improving overall performance.
Further observations show that using only the Forward part of HM-BiTCN outperforms the Backward part. This is closely related to the inherent unidirectionality of medical time-series signals such as EEG and ECG, where information typically propagates forward in time (e.g., neural signal transmission in EEG or atrial-to-ventricular activation in ECG). Such characteristics enable the Forward structure to capture key features and temporal evolution more effectively, yielding better performance. This finding not only deepens the understanding of medical signal processing but also provides insights for optimizing HM-BiTCN in related applications.
To evaluate the performance of our method on general time series, we follow the design of Medformer Wang et al. (2024a) and test it on two human activity recognition (HAR) datasets: FLAAP(13,123 samples, 10 classes) Kumar & Suresh (2022) and UCI-HAR(10,299 samples, 6 classes) Anguita et al. (2013).
Additionally, on the non-medical datasets FLAAP and UCI-HAR, we observed that integrating the bidirectional structure significantly improves performance. This indicates that in high-SNR scenarios, bidirectional modeling can more effectively capture both forward and backward feature information, enhancing overall model performance. In contrast, CIF provides relatively limited gains on these high-SNR datasets. This observation further highlights the design advantage of CIF: it is specifically tailored for low-SNR medical time series, explicitly fusing inter-channel physiological information to enhance signal quality and discriminative power, while its marginal benefit is smaller for low-noise non-medical data. Overall, these findings not only reveal the differential adaptability of model architectures under varying data characteristics but also underscore the unique value of CIF in complex medical scenarios.
Appendix E Further exploration of physiological structures
The parameters in CIF are explicit hyperparameters that can be directly set and adjusted based on experience. For example, Figure. 7 shows the corresponding locations of EEG channels on the human brain, we have adjusted the AFAVA dataset, which comprises 16 channels: C3, C4, F3, F4, F7, F8, Fp1, Fp2, O1, O2, P3, P4, T3, T4, T5, and T6. For the first six channels, we performed pairwise fusion as follows:
Here, the channels C3, F3, and F7 exhibit physiological symmetry with C4, F4, and F8, respectively. We denote this type of fusion as Physiological Symmetry Fusion (PSF). In contrast, the channel fusion previously described in our paper, which was adjusted via the hyperparameter , lacked physiological symmetry and is referred to as Random Fusion (RF).